OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8.

OpenStudy (anonymous):

a line parallel will have sam e slope of y = 1/2x - 8. so line slope=1/2 point (-2, 6) equation of the line is y-(6)=1/2(x-(-2)) can iu proceed

OpenStudy (anonymous):

yes